OpenStudy (anonymous):

Each of the numbers from 1 through 30 is written on a table tennis ball and placed in a wire cage. Each of the numbers 20 through 45 is written on a table tennis ball and placed in a different wire cage. One ball is chosen at random from each spinning cage. Which of the following best describes: P(each is a 25)

OpenStudy (anonymous):

Determine the probability that you draw a 25 from the first cage. Then determine the probability that you draw a 25 from the second cage. Then multiply these two probabilities together

OpenStudy (anonymous):

still dont get it

OpenStudy (anonymous):

Ok, how many balls are in the first cage? How many of them contain the number 25?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

Right, the first cage has 30. How many of them contain the number 25?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Ok, so the probability of getting a 25 from the first cage is 1/30 Now do the same thing for the second cage

OpenStudy (anonymous):

the second one has 1/25 right

OpenStudy (anonymous):

The second cage actually has 26 balls in it (the numbers 20 through 45) so the probability is 1/26

OpenStudy (anonymous):

Now, the probability that both of these events occur can be found by multipying them together: \[\frac{ 1 }{ 30 } \times \frac{ 1 }{ 26 }=\frac{ 1 }{ 30 \times 26 }\]

OpenStudy (anonymous):

1/780

OpenStudy (anonymous):

Right, that should be your answer